The cylinder head bolts on a car are to be tightened with a torque of 62.0 N⋅m. If a mechanic uses a wrench of length 20 cm, The perpendicular force must he exert on the end of the wrench to tighten a bolt correctly is _______


1
Expert's answer
2021-03-23T11:09:03-0400
"\tau=Frsin\theta,""F=\dfrac{\tau}{rsin\theta}=\dfrac{62\ N\cdot m}{0.2\ m\cdot sin90^{\circ}}=310\ N."